OpenStudy (anonymous):

Choose the graph which matches the following system of equations. 2x + y = –1 3x + 2y = –6

OpenStudy (anonymous):

we can firstly rearrange x-y=1 to be x=1 + y so y=x-1 then we sub y=x-1 into the second equation of 3x-2y=8 3x-2(x-1)=8 3x-2x+2=8 x=8-2 x=6 now sub x=6 back into either equations, lets choose y=x-1 so y=6-1 y=5 so the point of intersection here is (6,5)

OpenStudy (anonymous):

hope this worked